Hello,

I need help finding the volume of a cone with a radius of 10in and a height of 9in.

Thank you

Answer by checkley75(3666)

V=1/3B*H B=PI*R^2
V=1/3*PI*10^2*9
V=1/3*22/7*100*9
V=22/21*900
V=1.0476*900
V=942.84 IN^3